Pat is 28 years older than his son Patrick. In 8 years Pat will be twice as old as Patrick. How old are they now?
Patrick is ? years old
Pat is ? years old

Answer by robertb(5830)

Pat's present age = x + 28
Patrick's present age = x
Then 2(x + 8) = (x + 28) + 8
==> 2x + 16 = x + 36
==> 2x - x = 36 - 16
==> x = 20, the present age of Patrick, and
x + 28 = 48, the present age of Pat.